# Conditional normative reasoning as a fragment of HOL

### v3

@@ 1 Introduction
The paper is organized as follows. Section 5 recalls system E and its extensions. Section 3 shows the embedding of E in Isabelle/HOL. Section 4 studies the correspondence between the properties of the betterness relation and the axioms. Section 5 discusses the repugnant conclusion. Section 6 concludes. [^3] The theory files are available for downloading at http://logikey.org under sub-repository “/Deontic-Logics/cube-dll/” (files “DDLcube.thy”, “mere_addition_opt.thy”, “mere_addition_max.thy” and “mere_addition_lewis.thy”).
@@ 2.2 Semantics
Intuitively, quasi-transitivity demands that the strict part of the betterness relation be transitive. Acyclicity rules out strict betterness cycles. Suzumura consistency rules out cycles with at least one instance of strict betterness. It can readily be seen that acyclicity entails irreflexivity and asymmetry of $\succ$. Thus, the former may be interpreted as generalizing the latter two to a path of arbitrary length. Totality implies reflexivity. Given reflexivity and Ferrers, totality follows, and so the interval order condition can equivalently be defined by the pair “totality + Ferres”. Intuitively, the interval order condition makes room for the idea of non-transitive equal goodness relation due to discrimination thresholds. These are cases where $a\approx b$ and $b\approx c$ but $a\not\approx c$ (see Luce 1956).
@@
Lewis’s limit assumption is meant to rule out sets of worlds without a “limit” (viz. a best element). Its exact formulation varies among authors. It exists in (at least) the following four versions, where $\mathrm{best}\in\{\mathrm{max},\mathrm{opt}\}$
@@ 2.3 Systems
    $\Diamond\varphi\rightarrow(\bigcirc(\psi/\varphi)\rightarrow P(\psi/\varphi))$ (${}^{\star}$)
@@
We give an intuitive explanation for these axioms. COK is the conditional analog of the familiar distribution axiom K. Abs is the absoluteness axiom of Lewis, (1973), and reflects the fact that the ranking is not world-relative. Nec is the dyadic deontic counterpart of the familiar necessitation rule. Ext permits the replacement of necessarily equivalent formulas in the antecedent of deontic conditionals. Id is the deontic analog of the identity principle. Sh is named after (Shoham,, 1988, p. 77), who seems to have been the first to discuss it. One can see it as the deontic analog of one-half of the deduction theorem. ${}^{\star}$ is the conditional analog of the familiar D axiom. In its equivalent form, $\Diamond\varphi\rightarrow\neg(\bigcirc(\psi/\varphi)\wedge\bigcirc(\neg\psi/\varphi))$, this axiom rules out the possibility of conflicts between obligations arising in a context $\varphi$ that is possible. CM and DR correspond to the principle of cautious monotony and disjunctive rationality from the non-monotonic logic literature --Kraus et al., (1990). CM tells us that complying with an obligation does not modify the other obligations arising in the same context. DR tells us that if a disjunction of states of affairs triggers an obligation, then at least one disjunct triggers this obligation. Due to Spohn, (1975), Sp is best explained using the (more widely known) principle of rational monotony RM from non-monotonic logic--see Kraus et al., (1990). The two laws are inter-derivable in E. RM is obtained by replacing, in Sp, $\bigcirc(\psi\rightarrow\xi/\varphi)$ with $\bigcirc(\xi/\varphi)$, to read:
@@
Weak completeness establishes a match between theorems and validities only. This restriction is because the completeness proof appeals in an essential way to the assumption that models are finite--for more details, see (Parent,, 2024, §4).
@@ 2.4 Correspondences
limitedness ${}^{\star}$ ${}^{\star}$ Parent, (2015)
@@
To improve readability, we have used certain shortcuts, albeit with the potential drawback of simplifying the data. The lack of correspondence in the 1st, 2nd and 5th row (starting from the top, going downwards) is for the general case, when no constraint is put on $\succeq$. Thus, assuming one of reflexivity, totality or transitivity (under the max) does not add new validities. Similarly, the correspondence for limitedness is independent of any other properties (or axioms) in the background. The correspondence results for smoothness, transitivity (under the opt), transitivity+totality (under the max) and interval order assume ${}^{\star}$ and limitedness in the background. Quasi-transitivity, Suzumura consistency and acyclicity are known not to correspond to any formula in the general case, under the max rule--Parent, (2024). This holds whether or not limitedness or smoothness is assumed in the background. However, it is not known what happens under the opt rule. Therefore, we have put these three conditions aside.
@@ 3.2 Faithful embedding of system E
[…]
@@ 3.3 Properties
The other candidate weakenings of transitivity discussed earlier are also encoded in the theory file. For simplicity’s sake, we only give the example of quasi-transitivity and acyclicity. The encoding of the second is shown in Fig. 8. In Isabelle/HOL, the transitive closure of a relation can be defined in a few lines, shown in Fig. 9.
@@ 4.1 Max rule
  • If the relevant property is not assumed, counter-models for the corresponding axiom (${}^{\star}$, CM, DR and Sp) are found by Nitpick. This is Figs. 11, 13 and  14;
@@
Figure 11: ${}^{\star}$, CM and DR invalid in general
@@ 4.3 Inclusion
In Fig. 19, Sledgehammer shows the validity of the axioms of E holding independently of the properties assumed of the betterness relation. In Figs. 20 and 21, Sledgehammer confirms that the ${}^{\star}$ axiom and the Sp axiom call for totality and transitivity, respectively. Similarly, Fig. 22 shows that COK and CM call for both transitivity and totality. In all these cases, Sledgehammer fails to show the converse implications.
@@ 4.5 Assessment
As mentioned in Sect. 3.1, we found that the tools are very responsive in automatically proving (Sledgehammer), disproving (Nitpick), or showing consistency by providing a model (Nitpick). Not only an answer is returned, but also a justification for this answer. However, concerning this, we found that Nitpick fares better than Sledgehammer. When Sledgehammer has proved a theorem successfully, the list of the definitions, axioms and lemmas to be used is returned. But the derivation itself is not given. [^7] We tried to use Isar style reconstruction of proofs with Sledgehammer, but without much success for our examples. In principle one could look into the detailed proof output file of the external provers called by Sledgehammer, but this requires technical expertise. A simple example is given in Fig. 23. Line 170 tells us that quasi-transitivity follows from transitivity by assumption (“assms”) and using the definition of an asymmetric factor (“assfactor”). This is indeed how this would be shown by hand. However, a detailed argument is not given. By contrast, Nitpick always gives the full details of the model justifying its answer, and this one was always correct in our experiments. An example of such a model is given in Fig. 24, which we will discuss in a moment.
@@ 5.2 Computer-assisted experiments
Then $W$ contains an infinite increasing $\succ$-chain of $\neg A\wedge\neg A^{+}\!\!\wedge B$-worlds.
@@
Recall that quasi-transitivity entails acyclicity, which in turn entails irreflexivity and asymmetry of $\succ$ (Table 1). Assume that formulas (4)-(5)-(6) are satisfied; the argument primarily revolves around these. One may consider (3) and (7) as holding independently, without leading to a contradiction.
By (4), there is some $a_{1}$ such that $a_{1}\in\max(A\vee A^{+}\!)$ and $a_{1}\models A^{+}$. Hence, $a_{1}\models A^{+}\!\vee B$. By (5), $a_{1}\not\in\max(A^{+}\!\vee B)$, so there is some $a_{2}\succ a_{1}$ s.t. $a_{2}\models A^{+}\vee B$. By acyclicity (irreflexivity), $a_{1}$ and $a_{2}$ are distinct. Clearly, $a_{2}\models\neg A\wedge\neg A^{+}$. So $a_{2}\models B$, and hence by (6), $a_{2}\not\in\max(A\vee B)$. It follows that there is some $a_{3}\succ a_{2}$ s.t. $a_{3}\models A\vee B$. By acyclicity (irreflexivity and asymmetry), $a_{3}$ is other than $a_{2}$ and $a_{1}$. By quasi-transitivity, $a_{3}\succ a_{1}$. Since $a_{1}\in\max(A\vee A^{+}\!)$, $a_{3}\models\neg A\wedge\neg A^{+}$. Hence $a_{3}\models B$, and hence by (6) again, $a_{3}\not\in\max(A\vee B)$, and there is some $a_{4}$ s.t. $a_{4}\models A\vee B$ and $a_{4}\succ a_{3}$. By acyclicity, $a_{4}$ is other than $a_{1}$, $a_{2}$, $a_{3}$ and $a_{4}$. By quasi-transitivity, $a_{4}\succ a_{1}$, and so as before $a_{4}\models\neg A\wedge\neg A^{+}$. Reiterating the above argument indefinitely, one gets an infinite increasing $\succ$-chain of $\neg A\wedge\neg A^{+}\!\!\wedge B$-worlds, starting with $a_{2}$. ∎
@@ 6 Conclusion
Dr. X. Parent was funded in whole, or in part, by the Austrian Science Fund (FWF) [M3240 N, ANCoR project]. For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.

### v4

@@ 1 Introduction
The paper is organized as follows. Section 5 recalls system E and its extensions. Section 3 shows the embedding of E in Isabelle/HOL. Section 4 studies the correspondence between the properties of the betterness relation and the axioms. Section 5 discusses the repugnant conclusion. Section 6 concludes. [^3] The theory files are available for downloading at http://logikey.org under sub-repository “/Deontic-Logics/cube-dll/” (files “DDLcube.thy”, “mere_addition_opt.thy”, “mere_addition_max.thy” and “mere_addition_lewis.thy”). A corresponding (but slightly modified) Isabelle/HOL dataset is presented in Parent and Benzmüller, (2024).
@@ 2.2 Semantics
Intuitively, quasi-transitivity demands that the strict part of the betterness relation be transitive. Acyclicity rules out strict betterness cycles. Suzumura consistency rules out cycles with at least one instance of strict betterness. Acyclicity may be interpreted as generalizing asymmetry to a path of arbitrary length. Totality implies reflexivity. Given reflexivity and Ferrers, totality follows, and so the interval order condition can equivalently be defined by the pair “totality + Ferres”. Intuitively, the interval order condition permits instances where transitivity of equal goodness fails, due to discrimination thresholds. These are cases where $a\approx b$ and $b\approx c$ but $a\not\approx c$ (see Luce 1956).
@@
Lewis’s limit assumption is meant to rule out sets of worlds without a “limit” (viz. a best element). Its exact formulation varies among authors. It exists in (at least) the following four versions, where $\mathrm{best}\in\{\mathrm{max},\mathrm{opt}\}$:
@@ 2.3 Systems
    $\Diamond\varphi\rightarrow(\bigcirc(\psi/\varphi)\rightarrow P(\psi/\varphi))$ (D)
@@
We give an intuitive explanation for these axioms. COK is the conditional analog of the familiar distribution axiom K. Abs is the absoluteness axiom of Lewis, (1973), and reflects the fact that the ranking is not world-relative. Nec is the dyadic deontic counterpart of the familiar necessitation rule. Ext permits the replacement of necessarily equivalent formulas in the antecedent of deontic conditionals. Id is the deontic analog of the identity principle. Sh is named after (Shoham,, 1988, p. 77), who seems to have been the first to discuss it. One can see it as the deontic analog of one-half of the deduction theorem. D is the conditional analog of the familiar D axiom. In its equivalent form, $\Diamond\varphi\rightarrow\neg(\bigcirc(\psi/\varphi)\wedge\bigcirc(\neg\psi/\varphi))$, this axiom rules out the possibility of conflicts between obligations arising in a context $\varphi$ that is possible. CM and DR correspond to the principle of cautious monotony and disjunctive rationality from the non-monotonic logic literature --Kraus et al., (1990). CM tells us that complying with an obligation does not modify the other obligations arising in the same context. DR tells us that if a disjunction of states of affairs triggers an obligation, then at least one disjunct triggers this obligation. Due to Spohn, (1975), Sp is best explained using the (more widely known) principle of rational monotony RM from non-monotonic logic--see Kraus et al., (1990). The two laws are inter-derivable in E. RM is obtained by replacing, in Sp, $\bigcirc(\psi\rightarrow\xi/\varphi)$ with $\bigcirc(\xi/\varphi)$, to read:
@@
In part (i), (ii) and (iii) of Th. 2.1, and in Th. 2.2, soundness and completeness are taken in their strong sense. They establish a correspondence between the syntactic and semantic consequence relation while also accommodating a potentially infinite set of assumptions. To be more precise, the theorems are of the form: where $\Gamma$ is a set of formulas or assumptions, $\Gamma\vdash\varphi$ if and only if $\Gamma\models\varphi$. In part (iv) of Th. 2.1, soundness and completeness are taken in their weak sense: $\Gamma$ is required to be finite; this amounts to establishing a match between theorems and validities only. This restriction is because the completeness proof appeals in an essential way to the assumption that models are finite--for more details, see (Parent,, 2024, §4).
@@ 2.4 Correspondences
limitedness D D Parent, (2015)
@@
To improve readability, we have used certain shortcuts, albeit with the potential drawback of simplifying the data. The lack of correspondence in the 1st, 2nd and 5th row (starting from the top, going downwards) is for the general case, when no constraint is put on $\succeq$. Thus, assuming one of reflexivity, totality or transitivity (under the max) does not add new validities. Similarly, the correspondence for limitedness is independent of any other properties (or axioms) in the background. The correspondence results for smoothness, transitivity (under the opt), transitivity+totality (under the max) and interval order assume D and limitedness in the background. Quasi-transitivity, Suzumura consistency and acyclicity are known not to correspond to any formula in the general case, under the max rule--Parent, (2024). This holds whether or not limitedness or smoothness is assumed in the background. However, it is not known what happens under the opt rule. Therefore, we have put these three conditions aside.
@@ 3.2 Faithful embedding of system E
$\models^{\mathrm{PREF}}\varphi\text{ if and only if }\models^{\text{HOL}}\lfloor\varphi\rfloor$
@@ 3.3 Properties
The other candidate weakenings of transitivity discussed earlier are also encoded in the theory file. For simplicity’s sake, we only give the example of quasi-transitivity and acyclicity. The encoding of the second is shown in Fig. 8. In Isabelle/HOL, the transitive closure of a relation can be defined in a few lines, shown in Fig. 9. The encoding of quasi-transitivity is shown in Fig. 10.
@@ 4.1 Max rule
  • If the relevant property is not assumed, counter-models for the corresponding axiom (D, CM, DR and Sp) are found by Nitpick. This is Figs. 11, 13 and  14;
@@
Figure 11: D, CM and DR invalid in general
@@ 4.3 Inclusion
In Fig. 19, Sledgehammer shows the validity of the axioms of E holding independently of the properties assumed of the betterness relation. In Figs. 20 and 21, Sledgehammer confirms that the D axiom and the Sp axiom call for totality and transitivity, respectively. Similarly, Fig. 22 shows that COK and CM call for both transitivity and totality. In all these cases, Sledgehammer fails to show the converse implications.
@@ 4.5 Assessment
As mentioned in Sect. 3.1, we found that the tools are very responsive in automatically proving (Sledgehammer), disproving (Nitpick), or showing consistency by providing a model (Nitpick). Not only an answer is returned, but also a justification for this answer. However, concerning this, we found that Nitpick fares better than Sledgehammer. When Sledgehammer has proved a theorem successfully, the list of the definitions, axioms and lemmas to be used is returned. But the derivation itself is not given. [^7] We tried to use Isar style reconstruction of proofs with Sledgehammer, but without much success for our examples. In principle one could look into the detailed proof output file of the external provers called by Sledgehammer, but this requires technical expertise. A simple example is given in Fig. 23. Line 161 tells us that quasi-transitivity follows from transitivity by assumption (“assms”) and using the definition of an asymmetric factor (“assfactor”). This is indeed how this would be shown by hand. However, a detailed argument is not given. By contrast, Nitpick always gives the full details of the model justifying its answer, and this one was always correct in our experiments. An example of such a model is given in Fig. 24, which we will discuss in a moment.
@@ 5.2 Computer-assisted experiments
Then $W$ contains an infinite increasing $\succ$-chain of $\neg A\wedge\neg A^{+}\!\!\wedge B$-worlds (all distinct).
@@
I focus on the case where $\succeq$ is quasi-transitive. Recall that by definition $\succ$ is irreflexive and asymmetric, and that quasi-transitivity entails acyclicity (see Fig. 1). Assume that formulas (3)-(4)-(5) are satisfied; the argument primarily revolves around these. Equations (2) and (6) can also be assumed true without leading to a contradiction.
By (3), there is some $a_{1}$ such that $a_{1}\in\max(A\vee A^{+}\!)$ and $a_{1}\models A^{+}$. Hence, $a_{1}\models A^{+}\!\vee B$. By (4), $a_{1}\not\in\max(A^{+}\!\vee B)$, so there is some $a_{2}\succ a_{1}$ s.t. $a_{2}\models A^{+}\vee B$. By irreflexivity, $a_{1}$ and $a_{2}$ are distinct. Clearly, $a_{2}\models\neg A\wedge\neg A^{+}$. So $a_{2}\models B$, and hence by (5), $a_{2}\not\in\max(A\vee B)$. It follows that there is some $a_{3}\succ a_{2}$ s.t. $a_{3}\models A\vee B$. By irreflexivity and asymmetry, $a_{3}$ is other than $a_{2}$ and $a_{1}$. By quasi-transitivity, $a_{3}\succ a_{1}$. Since $a_{1}\in\max(A\vee A^{+}\!)$, $a_{3}\models\neg A\wedge\neg A^{+}$. Hence $a_{3}\models B$, and hence by (5) again, $a_{3}\not\in\max(A\vee B)$, and so there is some $a_{4}$ s.t. $a_{4}\models A\vee B$ and $a_{4}\succ a_{3}$. By acyclicity, $a_{4}$ is other than $a_{1}$, $a_{2}$, $a_{3}$ and $a_{4}$. By quasi-transitivity, $a_{4}\succ a_{1}$, and so as before $a_{4}\models\neg A\wedge\neg A^{+}$. Reiterating the above argument indefinitely, one gets an infinite increasing $\succ$-chain of $\neg A\wedge\neg A^{+}\!\!\wedge B$-worlds, starting with $a_{2}$.
The argument goes through if $\succeq$ is required to be transitive. This is because transitivity implies quasi-transitivity (see Fig. 1). ∎
Remark 1.
Note that Prop. 5.1 does not apply to to the interval order condition (even if this one implies quasi-transitivity as well). This is because (6) cannot simultaneously hold in a model where the interval order condition is satisfied. This can easily be verified. By (6), there is some $b_{1}\in\max(A^{+}\!\vee B)$ with $b_{1}\models B$. By (5), $b_{1}\not\in\max(A\vee B)$. So there is some $b_{2}\succ b_{1}$ with $b_{2}\models A\vee B$. Clearly, $b_{2}\models A\wedge\neg B\wedge\neg A^{+}\!$. We have $a_{1}\not\succeq a_{2}$ and $b_{1}\not\succeq b_{2}$. By Ferrers, $a_{1}\not\succeq b_{2}$ or $b_{1}\not\succeq a_{2}$. By totality, $b_{2}\succ a_{1}$ or $a_{2}\succ b_{1}$. The first contradicts the fact that $a_{1}\in\max(A\vee A^{+}\!)$, while the second contradicts the fact that $b_{1}\in\max(A^{+}\!\vee B)$. This is shown in Fig. 34, where the two cases are indicated by a “or”.
Figure 34: Adding (6)
@@ 6 Conclusion
Dr. X. Parent was funded in whole, or in part, by the Austrian Science Fund (FWF) [M3240 N, ANCoR project (doi: 10.55776/I2982), and 10.55776/I6372, LoDEX project]. For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.